# SPEED: Scalable, Precise, and Efficient Concept Erasure for Diffusion Models

### v1

@@ 5.2 On Multi-Concept Erasure
UCE 0.20 71.19 83.10 24.07 83.81 0.00 31.94 48.41 13.45 209.93 0.00 20.92 34.60 13.49 185.46
RECE 0.34 67.43 80.44 16.75 170.65 1.03 19.77 32.95 13.49 213.39 2.43 23.71 38.16 12.09 177.57
MACE 1.62 87.73 92.75 26.36 37.25 3.41 84.31 90.03 25.45 45.31 4.80 80.20 87.06 24.80 50.41
Table 3: Quantitative comparison of the multi-concept erasure in erasing 10, 50, and 100 celebrities. The best results are highlighted in bold. Our method is capable of erasing up to 100 celebrities at once with low $\text{Acc}_{e}$ (%) and preserving other non-target celebrities with less appearance alteration with high $\text{Acc}_{r}$ (%), resulting in the best overall erasure performance $H_{o}$ (shaded in pink).
NudeNet Detection Results on I2P MS-COCO
Armpits Belly Buttocks Feet Breasts (F) Genitalia (F) Breasts (M) Genitalia (M) Total ↓ CS ↑ FID ↓
SD v1.4 123 134 19 14 258 9 16 3 576 26.53 -
ConAbl 24 43 5 6 68 1 6 4 157 26.14 39.26
MACE 28 19 1 20 37 3 6 5 119 24.06 52.78
RECE 17 29 3 7 14 1 8 1 80 25.98 40.37
UCE 29 42 2 11 36 3 9 7 139 26.24 38.60
Ours 20 42 7 3 29 2 5 5 113 26.29 37.82
Table 4: Evaluation of implicit concept erasure on I2P benchmark. We report the number of nude body parts (F: Female, M: Male) detected by the NudeNet with threshold $=$ 0.6. The best and second-best results are marked in bold and underlined. (Left) Our method effectively removes nude content, even though nudity is not explicitly mentioned in prompts from I2P, achieving the second-best total count. (Right) Our method also consistently achieves superior prior preservation for non-target concepts to other methods on MS-COCO.
@@
Analysis and discussion. Table 3 showcases a notable improvement of our method on multi-concept erasure, particularly in prior preservation with the highest $\text{Acc}_{r}$. In comparison with the SOTA method, MACE [35], our method achieves superior prior preservation with better $\text{Acc}_{r}$, while maintaining comparable erasure efficacy, as reflected in similar $\text{Acc}_{e}$, resulting in the best overall erasure performance indicated by the highest $H_{o}$. Meanwhile, our method attains the lowest FID across all methods on MS-COCO. The other methods, UCE [18] and RECE [19], although achieving considerable balance in few-concept erasure, fail to maintain this balance as the number of target concepts increases as shown in Fig. 6, with catastrophic prior damage evidenced by MS-COCO as well. Notably, our method can erase up to 100 celebrities in 5 seconds, whereas MACE requires around 30 minutes (×$\times 350$ time). In real-world scenarios, this efficiency underscores our potential for the instant erasure of massive concepts.
5.3 On Implicit Concept Erasure
Evaluation setup. We further evaluate the erasure efficacy on implicit concepts, where the target concept does not explicitly appear in the text prompt. We conduct experiments on the Inappropriate Image Prompt (I2P) benchmark [51], which consists of various implicit inappropriate prompts involving violence, sexual content, and nudity. We follow the same setting in [19] to erase nudity $\to$ ‘ ’. Specifically, we generate images using all 4,703 text prompts in I2P and use NudeNet [6] to identify if the nude content is successfully erased with the threshold of 0.6. Additionally, we report the results on MS-COCO to demonstrate the prior preservation of general concepts.
Analysis and discussion. As shown in Table 4, our method can effectively erase the implicit concept, i.e., nudity, with the second-best number of detected nude body parts. The SOTA method, RECE [19], achieves the best total number by extending the erasure set with more target concepts, but this comes at the cost of sacrificing prior preservation on MS-COCO. In contrast, our method achieves the best prior preservation, demonstrating effective erasure while maintaining strong prior preservation, striking a favorable balance between erasure and preservation.
Figure 6: Quantitative comparison of multi-concept erasure in erasing different numbers of celebrities (celeb). The erased and preserved generations are highlighted with red and green boxes, respectively. Our method can precisely erase 100 celebrities while maintaining the proper generation of other non-target concepts.
@@ 5.4 Further Analysis
Duration comparison. Table 5 presents the duration comparison in erasing 1, 10, and 100 concepts across different methods. It is obvious that training-based methods necessitate significantly higher computational costs than editing-based ones. In contrast, our method achieves precise multi-concept erasure in a remarkably short time, demonstrating superior efficiency while maintaining erasure performance, as evidenced in Table 3.
Component ablation. In Table 6, we compare the individual impact of our components on prior preservation and draw the following conclusions: (1) Impact of IEC (Ablation 1 v.s. 2): IEC reduces the non-target FID and the MS-COCO FID, demonstrating its effectiveness by preserving invariant embeddings with equality constraints. (2) Impact of IPF (Ablation 2 v.s. 3): Incorporating IPF results in a significant improvement in both FIDs, underscoring its critical role in filtering out less-influenced concepts in the retain set under rank saturation. (3) Impact of DPA (Ablation 4 v.s. Ours): DPA improves RPA with directed noise and leads to a substantial improvement in non-target and MS-COCO FIDs, highlighting its advantage by introducing semantically similar concepts into the refined retain set. To conclude, the proposed three components (i.e., IEC, IPF, and DPA) improve the prior preservation from different perspectives and contribute to our method with the best prior preservation under null space constraints.
Training-based Editing-based
ConAbl MACE UCE RECE Ours
Data Prepare $n\times 1000t_{1}$ $n\times(8t_{1}+8t_{2})$ 0 0 0
1 concept $1\times 90$ 55.1 1.2 1.5 3.6
10 concepts $10\times 90$ 207.0 1.5 2.5 3.8
100 concepts $100\times 90$ 1750.9 2.1 11.0 5.0
Table 5: Duration comparison (s) in erasing multiple celebrities on one A100 GPU, where $n$ is the number of target concepts. During data preparation, ConAbl requires pre-sampling 1,000 images ($t_{1}$ per image) for each target concept while MACE needs 8 pre-sampled images along with 8 corresponding segmentation masks ($t_{2}$ per mask) using SAM [27].
@@
Table 6: Ablation study on the proposed components in erasing Van Gogh, with the non-target FID averaged over other four styles (i.e., Picasso, Monet, Paul Gauguin, Caravaggio) as in Table 2. Ablation 1 corresponds to the original objective from [14] in Eq. 3.2. The ablated components include: IEC (Invariant Equality Constraints), IPF (Influence-based Prior Filtering), RPA (Random Prior Augmentation), and DPA (Directed Prior Augmentation).
Figure 7: More applications across various T2I diffusion models. (a) We conduct both single and composite concept erasure for “Snoopy + Van Gogh” on two community T2I models [3, 4]. (b) Our method also enables model knowledge editing by specifying the anchor concept, such as “Wonder WomanWoman” and “SupermanBatman” on SDXL [44]. (c) Our method can seamlessly transfer to novel DiT-based T2I models (e.g., SDv3 [13]).
More applications on other T2I models. To validate the transferability of our method across versatile applications, we conduct further experiments on various T2I models with different weights and architectures, including: (1) Composite concept erasure on DreamShaper [3] and RealisticVision [4] from Fig 7 (a): Our method can precisely erase the target concept(s) while preserving other non-target elements within the prompt, such as the Van Gogh-style background (2nd column) and the Snoopy character (3rd column). (2) Knowledge editing on SDXL [44] from Fig 7 (b): The arbitrary nature of anchor concepts allows us to edit the pre-trained model knowledge. Herein, our method effectively edits the model knowledge while maintaining the overall layout and semantics of the generated images. (3) Instance erasure on SDv3 [13] from Fig 7 (c): To accommodate the diffusion transformer (DiT) [43] architecture in T2I models, we adapt our method to a DiT-based model, demonstrating a well-balanced trade-off between erasure (1st row) and preservation (2nd row) as well.
@@ 6 Conclusion
This paper introduced SPEED, a scalable, precise, and efficient concept erasure method for T2I diffusion models. It formulates concept erasure as a null-space constrained optimization problem, facilitating effective prior preservation along with precise erasure efficacy. Critically, SPEED overcomes the inefficacy of editing-based methods in multi-concept erasure while circumventing the prohibitive computational costs associated with training-based approaches. With our proposed Prior Knowledge Refinement involving three complementary techniques, SPEED not only ensures superior prior preservation but also achieves a 350×350$\times$350 × acceleration in multi-concept erasure, establishing itself as a scalable and practical solution for real-world applications.
@@ Appendix A Preliminaries
T2I diffusion models. T2I generation has seen significant advancements with diffusion models, particularly Latent Diffusion Models (LDMs) [49]. Unlike pixel-space diffusion, LDMs operate in the latent space of a pretrained autoencoder, reducing computational costs while maintaining high-quality synthesis. LDMs consist of a vector-quantized autoencoder [59, 12] and a diffusion model [11, 24, 55, 26, 57]. The autoencoder encodes an image $\boldsymbol{x}$ into a latent representation $\boldsymbol{z}=\mathcal{E}(\boldsymbol{x})$ and reconstructs it via $\boldsymbol{x}\approx\mathcal{D}(\boldsymbol{z})$. The diffusion model learns to generate latent codes through a denoising process. The training objective is given by [24, 49]:
$\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\boldsymbol{z}\sim\mathcal{E}(\boldsymbol{x}),\boldsymbol{c},\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{z}_{t},t,\boldsymbol{c})\right\|^{2}_{2}\right],$ (13)
where $\boldsymbol{z}_{t}$ is the noisy latent at timestep $t$, $\boldsymbol{\epsilon}$ is Gaussian noise, $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ is the denoising network, and $\boldsymbol{c}$ is conditioning information from text, class labels, or segmentation masks [49]. During inference, a latent $\boldsymbol{z}_{T}$ is sampled from a Gaussian prior and progressively denoised to obtain $\boldsymbol{z}_{0}$, which is then decoded into an image via $\boldsymbol{x}_{0}\approx\mathcal{D}(\boldsymbol{z}_{0})$.
Cross-attention mechanisms. Current T2I diffusion models usually leverage a generative framework to synthesize images conditioned on textual descriptions in latent space [49]. The conditioning mechanism is implemented through cross-attention (CA) layers. Specifically, textual descriptions are first tokenized into $n$ tokens and embedded into a sequence of vectors $\boldsymbol{e}\in\mathbb{R}^{d_{0}\times n}$ via a pre-trained CLIP model [45]. These text embeddings serve as the key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$ and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$ inputs using parametric projection matrices $\mathbf{W_{K}}\in\mathbb{R}^{d_{k}\times d_{0}}$ and $\mathbf{W_{V}}\in\mathbb{R}^{d_{v}\times d_{0}}$, while the intermediate image representations act as the query $\mathbf{Q}\in\mathbb{R}^{m\times d_{k}}$. The cross-attention mechanism is defined as:
@@ B.1 Deriving the Closed-Form Solution for UCE
$\mathop{\text{min}}\limits_{\boldsymbol{\Delta}}\left[\alpha\|(\mathbf{W}+\boldsymbol{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\beta\|\boldsymbol{\Delta}\mathbf{C}_{0}\|^{2}\right].$ (15)
To derive the closed-form solution, we begin by computing the gradient of the objective function with respect to $\boldsymbol{\Delta}$. The gradient is given by:
$\alpha\left(\mathbf{W}\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}+\boldsymbol{\Delta}\mathbf{C}_{1}\right)\mathbf{C}_{1}^{\top}+\beta\boldsymbol{\Delta}\mathbf{C}_{0}\mathbf{C}_{0}^{\top}=0.$ (16)
Solving the resulting equation yields the closed-form solution for $\boldsymbol{\Delta}_{\text{UCE}}$:
$\boldsymbol{\Delta}_{\text{UCE}}=\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right)^{-1}.$ (17)
@@
$\boldsymbol{\Delta}_{\text{UCE}}=\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I}\right)^{-1}.$ (18)
@@ B.2 Proof of the Lower Bound of $e_{0}$ for UCE
$e_{0}=\|\boldsymbol{\Delta}_{\text{UCE}}\mathbf{C}_{0}\|^{2}=\|\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I})^{-1}\mathbf{C}_{0}\|^{2}\geq c>0.$ (19)
@@
Define the projection of $\mathbf{B}$ onto the subspace spanned by the first $r$ right singular vectors (corresponding to the nonzero singular values) as
$\mathbf{B}_{\text{proj}}=\mathbf{V}_{r}^{\top}\mathbf{B},$ (27)
where $\mathbf{V}_{r}$ consists of the first $r$ columns of $\mathbf{V}$. Then, we can rewrite the expression as
@@
Next, we establish a lower bound for $\|\mathbf{B}_{\text{proj}}\|$. Since $\mathbf{B}$ is not entirely in the null space of $\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}$ (by assumption that the matrix is nonzero), there exists a constant $c_{1}>0$ such that
@@
B.3 Deriving the Closed-Form Solution for Ours

### v2

@@ 5.2 On Multi-Concept Erasure
$\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ Time ↓ CS ↑ FID ↓ $\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ Time ↓ CS ↑ FID ↓ $\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ Time ↓ CS ↑ FID ↓
SD v1.4 91.99 89.66 14.70 - 26.53 - 93.08 89.66 12.85 - 26.53 - 90.18 89.66 17.70 - 26.53 -
ConAbl 60.76 77.89 52.19 900 25.60 42.12 64.00 75.44 48.74 4,500 14.30 255.36 42.86 58.82 57.97 9,000 14.93 235.27
UCE 0.20 71.19 83.10 1.5 24.07 83.81 0.00 31.94 48.41 1.8 13.45 209.93 0.00 20.92 34.60 2.1 13.49 185.46
RECE 0.34 67.43 80.44 2.5 16.75 170.65 1.03 19.77 32.95 6.3 13.49 213.39 2.43 23.71 38.16 11.0 12.09 177.57
MACE 1.62 87.73 92.75 207 26.36 37.25 3.41 84.31 90.03 936 25.45 45.31 4.80 80.20 87.06 1736 24.80 50.41
Ours 1.81 89.09 93.42 3.8 26.47 30.02 3.46 88.48 92.34 4.2 26.46 39.23 5.87 85.54 89.63 5.0 26.22 44.97
@@
Analysis and discussion. Table 2 showcases a notable improvement of our method on multi-concept erasure, particularly in prior preservation with the highest $\text{Acc}_{r}$. In comparison with the SOTA method, MACE Lu et al. (2024a), our method achieves superior prior preservation with better $\text{Acc}_{r}$, while maintaining comparable erasure efficacy, as reflected in similar $\text{Acc}_{e}$, resulting in the best overall erasure performance indicated by the highest $H_{o}$. Meanwhile, our method attains the lowest FID across all methods on MS-COCO. The other methods, UCE Gandikota et al. (2024) and RECE Gong et al. (2025), although achieving considerable balance in few-concept erasure, fail to maintain this balance as the number of target concepts increases as shown in Fig. 5, with catastrophic prior damage evidenced by MS-COCO as well. Notably, our method can erase up to 100 celebrities in 5 seconds, whereas MACE requires around 30 minutes ($\times 350$ time). In real-world scenarios, this efficiency underscores our potential for the instant erasure of massive concepts.
Figure 6: More applications across various T2I diffusion models. (a) We conduct composite concept erasure for “Snoopy + Van Gogh” on DreamShaper Lykon (2023) (1st row) and RealisticVision SG161222 (2023) (2nd row). (b) Our method also enables model knowledge editing by specifying the anchor concept on SDXL Podell et al. (2023). (c) Our method can seamlessly transfer to novel DiT-based T2I models, e.g., SDv3 Esser et al. (2024).
@@ 5.3 Further Analysis
More applications on other T2I models. To validate the transferability of our method across versatile applications, we conduct further experiments on various T2I models with different weights and architectures, including: (1) Composite concept erasure on DreamShaper Lykon (2023) and RealisticVision SG161222 (2023) from Fig 6 (a): Our method can precisely erase the target concept(s) while preserving other non-target elements within the prompt, such as the Van Gogh-style background (2nd column) and the Snoopy character (3rd column). (2) Knowledge editing on SDXL Podell et al. (2023) from Fig 6 (b): The arbitrary nature of anchor concepts allows us to edit the pre-trained model knowledge. Herein, our method effectively edits the model knowledge while maintaining the overall layout and semantics of the generated images. (3) Instance erasure on SDv3 Esser et al. (2024) from Fig 6 (c): To accommodate the diffusion transformer (DiT) Peebles & Xie (2023) architecture in T2I models, we adapt our method to a DiT-based model, demonstrating a well-balanced trade-off between erasure (1st row) and preservation (2nd row) as well.
Table 3: Ablation study on proposed components in erasing Van Gogh, with the non-target FID averaged over the other four artistic styles from Table 1. Ablation 1 corresponds to the original objective from Fang et al. (2024) in Eq. 3. The ablated components include: IEC (Invariant Equality Constraints), IPF (Influence-based Prior Filtering), RPA (Random Prior Augmentation), and DPA (Directed Prior Augmentation).
@@
Component ablation. From Table 3, we compare the individual impact of our components on prior preservation and draw the following conclusions: (1) Impact of IEC (Ablation 1 v.s. 2): IEC reduces the non-target FID and the MS-COCO FID, demonstrating its effectiveness by preserving invariant embeddings with equality constraints. (2) Impact of IPF (Ablation 2 v.s. 3): Incorporating IPF results in a significant improvement in both FIDs, underscoring its critical role in filtering out less-influenced concepts in the retain set to mitigate semantic degradation. (3) Impact of DPA (Ablation 4 v.s. Ours): DPA improves RPA with directed noise and leads to a substantial improvement in non-target and MS-COCO FIDs, highlighting its advantage by introducing semantically similar concepts into the refined retain set. To conclude, the proposed three components (i.e., IEC, IPF, and DPA) improve the prior preservation from different perspectives and contribute to our method with the best prior preservation under null space constraints. More ablations are presented in Appx. D.5.
@@ 6 Conclusion
This paper introduced SPEED, a scalable, precise, and efficient concept erasure method for T2I diffusion models. It formulates concept erasure as a null-space constrained optimization problem, facilitating effective prior preservation along with precise erasure efficacy. Critically, SPEED overcomes the inefficacy of editing-based methods in multi-concept erasure while circumventing the prohibitive computational costs associated with training-based approaches. With our proposed Prior Knowledge Refinement involving three complementary techniques, SPEED not only ensures superior prior preservation but also achieves a $350\times$ acceleration in multi-concept erasure, establishing itself as a scalable and practical solution for real-world applications.
Ethics Statement
This work introduces a method for concept erasure in text-to-image diffusion models to address ethical concerns such as copyright infringement, privacy violations, and the generation of offensive content. By precisely removing specific target concepts while preserving the quality and semantics of non-target outputs, the proposed approach enhances the safety, reliability, and controllability of generative models. The method operates through parameter-space editing without requiring access to private data or involving human subjects, ensuring ethical integrity throughout the research process and promoting responsible deployment of generative AI technologies.
Reproducibility Statement
We have made extensive efforts to ensure reproducibility of our work. The proposed method, SPEED, is thoroughly described in the main paper (Secs. 3 and 4), with complete theoretical derivations provided in Appx. B. Implementation details, including experimental setup details and erasure configurations, are given in Appx. C. The experimental setups for all three erasure tasks (few-concept, multi-concept, and implicit concept erasure) are described in detail, with complete quantitative results and ablation studies reported in Sec. 5 and Appx. D.
@@ Appendix A Preliminaries
T2I diffusion models. T2I generation has seen significant advancements with diffusion models, particularly Latent Diffusion Models (LDMs) Rombach et al. (2022). Unlike pixel-space diffusion, LDMs operate in the latent space of a pretrained autoencoder, reducing computational costs while maintaining high-quality synthesis. LDMs consist of a vector-quantized autoencoder Van Den Oord et al. (2017); Esser et al. (2021) and a diffusion model Dhariwal & Nichol (2021); Ho et al. (2020); Sohl-Dickstein et al. (2015); Kingma et al. (2021); Song et al. (2020b). The autoencoder encodes an image $\bm{x}$ into a latent representation $\bm{z}=\mathcal{E}(\bm{x})$ and reconstructs it via $\bm{x}\approx\mathcal{D}(\bm{z})$. The diffusion model learns to generate latent codes through a denoising process. The training objective is given by Ho et al. (2020); Rombach et al. (2022):
$\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\bm{z}\sim\mathcal{E}(\bm{x}),\bm{c},\bm{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t,\bm{c})\right\|^{2}_{2}\right],$ (13)
where $\bm{z}_{t}$ is the noisy latent at timestep $t$, $\bm{\epsilon}$ is Gaussian noise, $\bm{\epsilon}_{\bm{\theta}}$ is the denoising network, and $\bm{c}$ is conditioning information from text, class labels, or segmentation masks Rombach et al. (2022). During inference, a latent $\bm{z}_{T}$ is sampled from a Gaussian prior and progressively denoised to obtain $\bm{z}_{0}$, which is then decoded into an image via $\bm{x}_{0}\approx\mathcal{D}(\bm{z}_{0})$.
Cross-attention mechanisms. Current T2I diffusion models usually leverage a generative framework to synthesize images conditioned on textual descriptions in the latent space Rombach et al. (2022). The conditioning mechanism is implemented through cross-attention (CA) layers. Specifically, textual descriptions are first tokenized into $n$ tokens and embedded into a sequence of vectors $\bm{e}\in\mathbb{R}^{d_{0}\times n}$ via a pre-trained CLIP model Radford et al. (2021). These text embeddings serve as the key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$ and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$ inputs using parametric projection matrices $\mathbf{W_{K}}\in\mathbb{R}^{d_{k}\times d_{0}}$ and $\mathbf{W_{V}}\in\mathbb{R}^{d_{v}\times d_{0}}$, while the intermediate image representations act as the query $\mathbf{Q}\in\mathbb{R}^{m\times d_{k}}$. The cross-attention mechanism is defined as:
@@ B.1 Deriving the Closed-Form Solution for UCE
$\mathop{\text{min}}\limits_{\bm{\Delta}}\left[\alpha\|(\mathbf{W}+\bm{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\beta\|\bm{\Delta}\mathbf{C}_{0}\|^{2}\right].$ (15)
To derive the closed-form solution, we begin by computing the gradient of the objective function with respect to $\bm{\Delta}$. The gradient is given by:
$\alpha\left(\mathbf{W}\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}+\bm{\Delta}\mathbf{C}_{1}\right)\mathbf{C}_{1}^{\top}+\beta\bm{\Delta}\mathbf{C}_{0}\mathbf{C}_{0}^{\top}=0.$ (16)
Solving the resulting equation yields the closed-form solution for $\bm{\Delta}_{\text{UCE}}$:
$\bm{\Delta}_{\text{UCE}}=\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right)^{-1}.$ (17)
@@
$\bm{\Delta}_{\text{UCE}}=\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I}\right)^{-1}.$ (18)
@@ B.2 Proof of the Lower Bound of $e_{0}$ for UCE
$e_{0}=\|\bm{\Delta}_{\text{UCE}}\mathbf{C}_{0}\|^{2}=\|\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I})^{-1}\mathbf{C}_{0}\|^{2}\geq c>0.$ (19)
@@
Define the projection of $\mathbf{B}$ onto the subspace spanned by the right singular vectors as
$\mathbf{B}_{\text{proj}}=\mathbf{V}^{\top}\mathbf{B}.$ (27)
Then, we can rewrite the expression as
@@
Next, we establish a lower bound for $\|\mathbf{B}_{\text{proj}}\|$. Given that $\mathbf{V}$ is composed of right singular vectors, there exists a smallest non-zero singular value $c_{1}$ such that:
@@
B.3 Deriving the Closed-Form Solution for SPEED